# Continuous Test-time Domain Adaptation for Efficient Fault Detection under Evolving Operating Conditions

### v1

@@ 4.3 Scoring and Fault Detection
During test time, we compute the fault label $y\in[0,1]$ based on the reconstruction result. 00 denotes a healthy sample while $1$ indicates a faulty sample. Given the $i_{th}$ data sample $X_{i}=[x_{i}^{1},....x_{i}^{k}]$, we compute its relative residual:
@@
given its predicted reconstruction result $\hat{X_{i}}$. $k$ indicates the input dimension. $\bar{X}_{t\_training}$ represents the mean value of target data samples for training (including validation data), which helps scale the residual values. The anomaly score $s_{i}$ is calculated by integrating the scaled residual values across all sensors:
@@
To avoid false detection by outliers with extremely large residuals, the computed anomaly score is smoothed within a defined window length $l$:
$s_{i\_smooth}=min\sum_{q=0}^{l-1}s_{i+q}$ (4)
@@ 5.1 Implementation details
Baselines We evaluate the performance of our proposed method against two established domain adaptation techniques: Adaptive Batch Normalization (AdaBN) [41] and Maximum Mean Discrepancy (MMD) [42]. Both methods are unsupervised and do not require labels in the source or target domains, making them suitable for comparison with our approach. For MMD, we align the feature distributions of the source and target domains by minimizing the discrepancy during training, using the same target domain data as in our proposed approach. Methods such as Domain-Adversarial Neural Networks (DANN) [43], which require labeled source data during adaptaion, are not included in our evaluation, as they operate under different assumptions and are therefore not directly comparable. Additionally, we include a baseline model trained exclusively on source domain data without any adaptation, serving as a reference to quantify the improvement achieved by domain adaptation strategies.
@@ 5.2 Case study 1: Cranfield three-phase flow facility
Dataset The Three-Phase Flow Facility at Cranfield University is designed to provide a controlled and measurable flow of water, oil, and air within a pressurized system [44]. Table 1 shows the monitored process variables For model training, three datasets (T1, T2, and T3) were collected under normal operating conditions, with varying air and water flow rates to ensure data diversity. In addition to the normal operation datasets, data were collected under faulty conditions, where six distinct faults were intentionally introduced to the system. These faults were selected to simulate real-world malfunctions. The faults were introduced after a period of stable normal operation for the investigation of changes in health indicators. Whenever feasible, the faults were applied gradually, allowing observation of how increasing fault severity impacted the system behavior. Once the fault reached a predefined severity level, it was removed, and the system was returned to normal operating conditions.
Table 2 summarizes the fault cases introduced in the Three-Phase Flow Facility, providing details on the operation condition, fault duration, and the start and end times for each injected fault. Six fault types were included: air line blockage, water line blockage, top separator input blockage, open direct bypass, slugging conditions, and pressurization of the 2” line. In all cases, faults were introduced within defined time windows while the system operated under either steady or changing conditions. The duration and timing of each fault event varied across the datasets. For the slugging conditions, faults were injected multiple times within a single sub-dataset.
@@
Fault case Operation condition Duration Fault start (s) Fault end (s)
Air line blockage changing 5811 1566 5181
steady 4467 657 3777
steady 4321 691 3691
Water line blockage changing 9192 2244 6616
steady 3496 476 2656
steady 3421 331 2467
Top separator input blockage changing 9090 1136 8352
steady 6272 333 5871
steady 10764 596 9566
Open direct bypass changing 7208 953 6294
steady 4451 851 3851
steady 3661 241 3241
Slugging conditions changing 2541 686 1772
1772 2253
changing 10608 1633 2955
7031 7553
8057 10608
Pressurization of the 2” line changing 2800 1723 2800
changing 4830 1037 4830
@@
Implementation details Our objective is to enable robust fault detection with limited data availability. A source model is first pretrained with abundant source data, and then adapted to target operation conditions that have limited colleted data samples. In the Cranfield case study, the training dataset (T1, T2 and T3) collected under varying normal operating conditions is selected as the source domain. Each faulty case study is treated as a separate target domain, where only a limited number of training samples are available. Specifically, for each fault case, we use 2500 normal samples for training and evaluate the model on the remaining data. During testing, the training data from each target domain is utilized to compute the anomaly detection threshold $\bar{r}_{t_{t}raining}$. The detection sensitivity parameter is set to $\alpha=1.0$ for all case studies in this experiment.
Fault detection result Table 3 summarizes the fault detection performance across six fault cases, comparing our proposed TARD method with the baseline, AdaBN, and MMD adaptation methods. Overall, the proposed TARD method consistently outperforms the other methods across different faulty cases, highlighting its effectiveness and robustness in fault detection, particularly under limited training data conditions. In challenging fault cases, such as airline blockage and open direct bypass, where both the baseline and other adaptation methods perform relatively poorly (F1 scores around or below 0.56), TARD achieves notable improvements, reaching F1 scores of 0.70 and 0.69, respectively. These results indicate that TARD is more robust in detecting subtle or less distinguishable fault patterns. In fault cases which are relatively easy to detect, such as slugging conditions and pressurization of the 2” line, most methods already achieve high performance, while TARD still maintains the highest or comparable metrics, with F1 scores of 0.92 and 1.00, showing consistent improvement across various adaptation scenarios.
@@ 5.3 Case study 2: PRONTO heterogeneous benchmark
Dataset The Pronto heterogeneous benchmark dataset is an industrial-scale multiphase flow facility for the investigation of the dynamics and control of multiphase flows. Table 5 shows the monitored process variables. A total of 20 distinct operating conditions are covered in the dataset, characterized by various combinations of air and water flow rates. In addition to normal operation, the dataset incorporates three types of manually induced faults, including air leakage, air blockage, and diverted flow. Air leakage involves a gradual loss of air supply, leading to cyclic flow behavior and eventually a transition to a liquid-only regime. Air blockage restricts the air supply without reducing system pressure, resulting in milder forms of slugging. Diverted flow reroutes the flow, altering the regime, particularly at the riser section. Each fault type is applied at two different operating points, with the severity adjusted incrementally. Each of these faults was introduced at two different operating points by incrementally adjusting valve openings, allowing the fault severity to be progressively varied. The dataset includes precise timestamps, process variable measurements, and supplementary data such as alarm logs, high-frequency sensor readings, and visual recordings, all aligned with the induced faults. For comprehensive fault descriptions, please refer to [45].
@@ 6 Conclusions
Despite its strong performance, we see potential directions for improvements in the current method. First, our adaptive module uniformly adapts to all the incoming batches without quantifying the extent of the domain shifts. We hypothesize that the performance could be enhanced by re-training this module once a significant domain shift is detected. Second, the thresholding parameter $\alpha$, which determines the sensitivity of fault detection, is currently set manually based on empiracle experience. An automatic adjustment of this parameter, taking into account both distribution shifts and operational requirements, could optimize the trade-off between minimizing false alarms and attaining prompt fault detection.

### v2

@@ 4.3 Scoring and Fault Detection
During test time, we compute the fault label $y\in[0,1]$ based on the reconstruction result. 0 denotes a healthy sample while $1$ indicates a faulty sample. Given the $i_{th}$ data sample $X_{i}=[x_{i}^{1},....x_{i}^{k}]$, we compute its relative residual:
@@
given its predicted reconstruction result $\hat{X_{i}}$. $k$ indicates the input dimension. $\bar{X}_{t\_training}$ represents the mean value of target data samples for training (including validation data), which helps scale the residual values. The anomaly score $s_{i}$ for each indidual data sample is calculated by integrating the scaled residual values across all of its $k$ sensor measurements:
@@
To avoid false detection by outliers with extremely large residuals, the computed anomaly score is smoothed within a defined window length $l$ across the time dimension:
$s_{i\_smooth}=mean\sum_{q=0}^{l-1}s_{i+q}$ (4)
@@ 5.1 Implementation details
Training details The adaptation training for our TARD module and the baseline methods was conducted with a consistent set of hyperparameters. We used the Adam optimizer with a learning rate of 1e-5. All models were trained with a batch size of 128. The basedline autoencoder was trained for 500 epochs. For TARD, the adaptive module ( $h_{\phi}$) was trained for 50 epochs, as its small size led to rapid convergence. During this phase, all batch normalization layers in the main autoencoder ($f_{\theta}$) were frozen. The AdaBN layer within our adaptive module, as well as in the AdaBN baseline, updates its running statistics on a per-batch basis during both adaptation training and subsequent testing.
Baselines We evaluate the performance of our proposed method against two established domain adaptation techniques: Adaptive Batch Normalization (AdaBN) [42] and Maximum Mean Discrepancy (MMD) [43]. These baselines were specifically selected to ensure a fair and relevant comparison under the strict constraints of our problem setting: complete absence of labeled fault data, limited normal target data, and no access to source data during online adaptation. Both methods are unsupervised and do not require labels in the source or target domains, making them suitable for comparison with our approach. Furthermore, both they represent widely-adopted approaches for unsupervised domain adaptation—AdaBN for feature-level normalization strategies and MMD for explicit distribution alignment—making them appropriate benchmarks for evaluating our proposed method’s. For MMD, we align the feature distributions of the source and target domains by minimizing the discrepancy during training, using the same target domain data as in our proposed approach. Methods such as Domain-Adversarial Neural Networks (DANN) [37], which require labeled source data during adaptaion, are not included in our evaluation, as they operate under different assumptions and are therefore not directly comparable. Similarly, we exclude methods that require simultaneous access to source and target data during training due to practical deployment constraints where source data may be unavailable. Additionally, we include a baseline model trained exclusively on source domain data without any adaptation, serving as a reference to quantify the improvement achieved by domain adaptation strategies.
@@ 5.2 Case study 1: Cranfield three-phase flow facility
Dataset The Three-Phase Flow Facility at Cranfield University is a laboratory-scale setup designed to provide a controlled and measurable flow of water, oil, and air within a pressurized system [44]. The facility consists of key industrial components, including multiple pipelines, a 10.5-meter-high riser, a two-phase separator, and a larger three-phase separator at ground level, along with compressors, pumps, and storage tanks. The entire process is managed by a SCADA system. The system is monitored by 24 different process variables, which include pressures, flow rates, levels, densities, temperatures, and valve positions. Table 1 shows the monitored process variables. The two primary inputs used to control the operating conditions are the air and water flow rate set points. For model training, three datasets (T1, T2, and T3) were collected under normal operating conditions, with varying air and water flow rates to ensure data diversity. In addition to the normal operation datasets, data were collected under faulty conditions, where six distinct faults were intentionally introduced to the system. These include air and water line blockages, a top separator input blockage, an open direct bypass to simulate a leakage, slugging conditions to create an unstable flow regime, and the pressurization of the 2” line to simulate an incorrect operational procedure. These faults were selected to simulate real-world malfunctions. The faults were introduced after a period of stable normal operation for the investigation of changes in health indicators. Whenever feasible, the faults were applied gradually, allowing observation of how increasing fault severity impacted the system behavior. Once the fault reached a predefined severity level, it was removed, and the system was returned to normal operating conditions.
Table 2 summarizes the fault cases introduced in the Three-Phase Flow Facility, providing details on the operation condition, time span, the ratio of faulty samples among all data samples. In this study, all the data was captured at a sampling rate of 1 Hz. Six fault types were included: air line blockage, water line blockage, top separator input blockage, open direct bypass, slugging conditions, and pressurization of the 2” line. In all cases, faults were introduced within defined time windows while the system operated under either steady or changing conditions. The duration and timing of each fault event varied across the datasets. For the slugging conditions, faults were injected multiple times within a single sub-dataset.
[…]
Figure 3: Visualization of the first two principle components of PCA analysis of the Cranfield three-phase flow facility dataset, including the normal training dataset and the testing dataset for each faulty condition.
@@
Fault case Operation condition Duration (s) Fault ratio
Air line blockage changing 5811 0.62
steady 4467 0.70
steady 4321 0.69
Water line blockage changing 9192 0.48
steady 3496 0.62
steady 3421 0.62
Top separator input blockage changing $90$ 0.79
steady 6272 0.88
steady 10764 0.83
Open direct bypass changing 7208 0.74
steady 4451 0.67
steady 3661 0.82
Slugging conditions changing 2541 0.43
0.19
changing 10608 0.12
0.05
0.24
Pressurization of the 2” line changing 2800 0.38
changing 4830 0.79
@@
Implementation details Our objective is to enable robust fault detection with limited data availability. A source model is first pretrained with abundant source data, and then adapted to target operation conditions that have limited collected data samples. In the Cranfield case study, the training dataset (T1, T2 and T3) collected under varying normal operating conditions is selected as the source domain. Each faulty case study is treated as a separate target domain, where only a limited number of training samples are available. Specifically, for each fault case, we use 2500 normal samples for training and evaluate the model on the remaining data. During testing, the training data from each target domain is utilized to compute the anomaly detection threshold $\bar{r}_{t_{t}raining}$. The detection sensitivity parameter is set to $\alpha=1.0$ for all case studies in this experiment.
Fault detection result Table 3 summarizes the fault detection performance across six fault cases, comparing our proposed TARD method with the baseline, AdaBN, and MMD adaptation methods. Overall, the proposed TARD method consistently outperforms the other methods across different faulty cases, highlighting its effectiveness and robustness in fault detection, particularly under limited training data conditions. While the baseline and other adaptation methods perform relatively poorly in challenging cases such as airline blockage and open direct bypass (F1 scores around or below 0.56), TARD achieves notable improvements, reaching F1 scores of 0.70 and 0.69, respectively. These results indicate that TARD is more robust in detecting subtle or less distinguishable fault patterns. In fault cases which are relatively easy to detect, such as slugging conditions and pressurization of the 2” line, most methods already achieve high performance, while TARD still maintains the highest or comparable metrics, with F1 scores of 0.92 and 1.00, demonstrating consistent improvement across various adaptation scenarios.
The superior performance of TARD can be attributed to its dedicated adaptation strategy. Unlike AdaBN, which adjusts all model parameters and can overfit to batch statistics of potential faulty samples, and MMD, which aligns global distributions and may miss local shifts online, TARD’s separate adaptive module focused on control variables allows it to precisely compensate for operational changes without distorting the learned representation of normal sensor behavior. This targeted adaptation is particularly effective in reducing false alarms caused by distribution shifts.
@@
A key observation in Figure 4 is the apparent delay between the start of the reported ground truth (GT) fault period and the anomaly detected by TARD. This is an expected result that reflects the physical nature of the fault. The ’Top separator input blockage’ is a gradually developing fault. The GT period marks the initiation of the blockage, but it takes time for its effects to propagate and manifest as a significant deviation in the sensor readings. The initial, subtle changes are characteristic of an incipient fault, which may not be distinguishable from normal process noise. Our method, with its threshold set for robustness, correctly identifies the fault once its signature becomes statistically significant and persistent, thus ensuring reliable detection while avoiding false alarms from minor, ambiguous fluctuations.
@@
To further discuss the uncertainty in our fault detection results, we conducted an uncertainty quantification (UQ) experiment for this Cranfield Top separator input blockage case study, where we implemented a Deep Ensemble for estimating the uncertainty in the anomaly score. We trained 10 independent TARD models with identical architectures but different random weight initializations. During inference, each test batch is processed by all 10 models in the ensemble. This yields a distribution of 10 different anomaly scores for each data point. The mean of this distribution is used as the final, more robust anomaly score, while the standard deviation serves as a direct measure of the models’ disagreement, or uncertainty.
Figure 5 displays the mean anomaly score from the deep ensemble, with the shaded blue band representing the model uncertainty (± one standard deviation). During the detected fault period, the uncertainty band remains consistently narrow, with the standard deviation averaging approximately 0.8. This low level of disagreement among the ensemble models signifies a high-confidence consensus. This result confirms that the detection is not only effective but also highly robust.
Figure 5: Uncertainty estimation for the Cranfield Top separator input blockage case. The gray regions indicate the reported fault periods for each dataset subset and the red regions indicate the predicted anomaly region. The red line denotes the computed anomaly detection threshold $\alpha*\bar{r}_{t\_training}$. The shaded blue confidence interval represents mean anomaly score ± one standard deviation
@@ 5.3 Case study 2: PRONTO heterogeneous benchmark
Dataset The Pronto heterogeneous benchmark dataset is an industrial-scale high-pressure multiphase flow facility for the investigation of the dynamics and control of multiphase flows. The process for this study involved mixing water and air, which then flowed through a horizontal pipeline to a 2-inch riser with an S-shaped section before being separated; the oil phase was not used. The facility is heavily instrumented with standard pressure, flow, temperature, and density sensors, as well as high-frequency pressure and ultrasonic sensors. Table 6 shows the monitored process variables. A total of 20 distinct operating conditions are covered in the dataset, characterized by various combinations of air and water flow rates. Table 5. summarizes the fault cases introduced in the dataset, providing details on the operation condition, time span, the ratio of faulty samples among all data samples. In this study, all the data was captured at a sampling rate of 1 Hz. The dataset incorporates three types of manually induced faults, including air leakage, air blockage, and diverted flow. Air leakage involves a gradual loss of air supply, leading to cyclic flow behavior and eventually a transition to a liquid-only regime. Air blockage restricts the air supply without reducing system pressure, resulting in milder forms of slugging. Diverted flow reroutes the flow, altering the regime, particularly at the riser section. Each fault type is applied at two different operating points, with the severity adjusted incrementally. Each of these faults was introduced at two different operating points by incrementally adjusting valve openings, allowing the fault severity to be progressively varied. The dataset includes precise timestamps, process variable measurements, and supplementary data such as alarm logs, high-frequency sensor readings, and visual recordings, all aligned with the induced faults. For comprehensive fault descriptions, please refer to [45].
Fault case Duration (s) Fault ratio
Air leakage 6009 0.21
Air blockage 5947 0.20
Diverted flow 7143 0.33
Table 5: Summary of introduced faults of PRONTO dataset
We visualize the first two principal components of PCA of the PRONTO dataset, which account for $92\%$ of the data’s variability. The plot clearly illustrates the domain shift between the ’Normal’ source data (blue) and the three fault conditions. The separation confirms that the operating points where faults were introduced are distinct from the source domain, validating the need for an adaptation method like TARD. Furthermore, the visualization reveals the complex, non-linear structure of the fault data. While ’Air blockage’ (green) forms relatively tight clusters, ’Air leakage’ (orange) and especially ’Diverted flow’ (red) are widely distributed, highlighting the challenging nature of detecting these conditions. The significant overlap between fault types underscores the focus of our work on fault detection (distinguishing normal from any fault) in complex systems.
Figure 6: Visualization of the first two principle components of PCA analysis of the PRONTO heterogeneous dataset, including the normal training dataset and the testing dataset for each faulty condition.
@@
The results on the PRONTO dataset further validate TARD’s effectiveness. The significant improvement on the ”diverted flow” fault, which is particularly subtle, underscores the advantage of TARD’s targeted adaptation. By focusing adaptation on the control variables, TARD avoids overfitting to the overall data distribution and maintains sensitivity to nuanced fault signatures that other methods miss.
Fault Case Baseline AdaBN MMD TARD (ours)
Acc F1 AUC Acc F1 AUC Acc F1 AUC Acc F1 AUC
Air leakage 0.81 0.62 0.78 0.81 0.36 0.63 0.74 0.51 0.74 0.87 0.76 0.92
Air blockage 0.97 0.93 0.96 0.94 0.88 0.96 0.99 0.96 0.97 0.98 0.94 0.98
5.4 Case study 3: High-dimensional synthetic dataset
To validate the scalability and performance in high-dimensional settings, we generated a synthetic benchmark dataset. The simulated features include control variables $w$, sensor measurements $x$, and normalized time index $t\in[0,1]$ to simulate a gradual temporal shift. The input features consist of independent control variables $w$. A high-dimensional set of sensor measurements $x$ was generated as non-linear functions of these control variables. To simulate a realistic industrial environment, the sensor responses were designed with intricate interdependencies, including:
  • 1.
    Temporal Domain Drift: We incorporate the evolution over time by linearly drifting some of the sensor measurements linearly over time. This simulates real-world phenomena such as equipment wear and tear, gradual calibration shifts, or changing environmental conditions, ensuring the ”normal” state is a moving target.
  • 2.
    Redundancy and Correlation: To simulate the interdependency common in industrial systems, a subset of sensors was defined as linear combinations of others. These redundant sensors inherently inherit the drift of their source signals.
  • 3.
    Stochastic Noise: Gaussian noise was added to each sensor signal to mimic real-world measurement imprecision.
We generate datasets for two sets of experiments: one with $100$ sensor measurements and $10$ control variables, and another in a higher-dimensional setting with $1000$ sensor measurements and $100$ control variables. For each setting, we generate normal data of 2,000 timesteps for pre-training. To simulate abnormal behavior, we synthesize a test set of 5,000 data samples and insert two contiguous anomaly periods covering approximately 10% of the 5,000 timesteps. During each anomaly window, a random subset of state features is shifted by a constant offset. All timesteps within these windows are labeled anomalous ($\texttt{label}=1$), while the rest are normal ($\texttt{label}=0$). The final dataset therefore has shape $5000\times 111$, consisting of 100 state features, 10 control features, and 1 binary label column.
Table 9 summarizes the detection results and computational requirements for both scenarios.In both settings, TARD significantly outperforms the baseline. Although the performance is challenged in the extreme case of 1000 sensor measurements, it still provides a substantial improvement over the baseline, which struggles significantly in this high-dimensional space. The results also demonstrate that TARD’s adaptation process is highly efficient. The inference latency remains well within the requirements for real-time monitoring, confirming that TARD efficiently scales for high-dimensional data.
Sensor Measurements Method Detection Result Computational Efficiency
Acc F1 Training Time (s) Inference Time (ms) Number of Model Parameters
[…]
Table 9: Experimental results on high-dimensional synthetic dataset.
The TARD framework is well-suited for systems with many sensors and complex dependencies, as confirmed by these results. The core of the model is a reconstruction-based autoencoder that is pre-trained offline on the source dataset. This autoencoder, composed of fully-connected layers, is chosen to effectively learn a low-dimensional representation of the non-linear relationships hidden within the high-dimensional sensor data. Crucially, the online adaptation is handled by a separate, lightweight adaptive module. This module is intentionally kept simple, consisting of only two fully connected layers. Its size and complexity scale only with the number of control variables, not the total number of sensors, as it exclusively processes control variables as its input.
The computational and memory loads are intentionally split between offline and online phases to ensure real-time feasibility. The primary memory requirement is for storing the weights of the pre-trained autoencoder, which is a one-time cost determined before deployment. The most computationally intensive phase is the initial training of the autoencoder on the source data, but since this is performed only once, it does not impact the model’s operational performance. At test time, the process is highly efficient. It involves a single forward pass through the frozen autoencoder and the very small adaptive module. The adaptation is continuous during inference, with an AdaBN layer in the adaptive module updating its statistics based on each test batch. This lean online process makes TARD suitable for real-time fault detection.
@@ 6 Conclusions
Despite its strong performance, we have identified key directions for future enhancement to bolster TARD’s autonomy and long-term reliability in industrial settings. A primary area for improvement is the development of a robust safeguard mechanism to handle major, permanent system changes. This would involve continuously quantifying the domain shift in control variables and using a threshold to trigger a targeted retraining of the adaptive module ($h_{\phi}$) when a significant shift is detected. This ensures the model remains synchronized with the system’s evolving baseline without requiring a full reset. Furthermore, developing a method for the automatic adjustment of the fault detection sensitivity parameter ($\alpha$) would create a more adaptive and self-sufficient system, optimizing the trade-off between minimizing false alarms and ensuring prompt fault detection in dynamic environments. Finally, the fault analysis can be enhanced by examining the temporal dynamics of the anomaly scores themselves. While our current approach smooths residuals to ensure robustness against transient sensor noise, this trades temporal detail for reliability. A powerful next step is to introduce a secondary analysis module, such as a dedicated time-series model, to specifically detect complex patterns like slow-growing trends or cyclical behaviors within the residual sequence.